# On the degrees of freedom of spatially covariant vector field theory

### v1

@@ V.2.1 Case 1
We need to calculate the Poisson bracket $\ddot{\phi}_{1}(\vec{x}),H_{\mathrm{C}}]$ with $\ddot{\phi}_{1}=\mathcal{C}_{k}^{*(0,i)}\Pi^{k}+\mathcal{D}^{*(0,i)}$. We find
$[\ddot{\phi}_{1}(\vec{x}),H_{\mathrm{C}}]$ $=\Theta_{m}^{(0,i)}\Pi^{m}+\Delta^{(0,i)},$ (88)
@@
$\Theta^{m(0,i)}=$ $\mathcal{A}\mathcal{K}\Bigg[\frac{1}{\mathcal{A}}\frac{\partial\mathcal{D}^{*(0,i)}}{\partial A_{m}}-\partial_{k}\left(\frac{\partial^{2}\mathcal{V}^{(0,i)}}{\partial A_{m}\partial A_{k}}\right)+\partial^{m}\left(\frac{\partial^{2}f_{3,2}}{\partial A_{l}\partial A_{k}}\right)\partial_{k}A_{l}$
@@
which indicates the emergence of a first-class constraint and thus the enhancement of the gauge symmetry. The resulting theory possesses one first-class constraint and two second-class constraints, which thus belongs to “type-I” theories. Comparing with the case in Sec. V.1.2, if the conditions discussed in V.1.2 are satisfied, the Poisson brackets $\phi_{1},\ddot{\phi}_{1}]\approx 0$ and $\ddot{\phi}_{1},H_{\mathrm{C}}]\approx 0$ will automatically be satisfied, but not vice versa. In other words, the case 2 of branch 1 discussed in V.1.2 is not independent in the sense that it can be viewed as the sub-case discussed here, i.e., as the special solution in case 1 of branch 2.
We then turn to the general case, where $\ddot{\phi}_{1}(\vec{x}),H_{\mathrm{C}}]$ is not vanishing weakly. This leads to a new quaternary constraint, yielding the Dirac matrix
@@
$\tilde{\mathcal{F}}_{k}^{*}=-f_{4,2}\Theta_{k}^{(0,i)}$ (94)
@@
We refer to theories in this case “type-II” theories. It is clear that this is a new situation different from the previous ones.
@@
where $f_{2}$ is a nonlinear polynomial of $\bar{X}$. For $\mathcal{L}_{3}$, we find that $f_{3,1}$ cannot depend on $\bar{X}$ but can be a general polynomial of $A_{0}$, while $f_{3,2}$ must be nonlinear in $\bar{X}$. Thus we can write
@@
$h_{4,1}$ $=\alpha_{2},\quad f_{4,2}=\beta_{2}\quad,h_{4,3}=\gamma_{2},$ (99)
@@ V.2.2 Case 2
Recalling that $\dot{\phi}_{1}=\mathcal{A}\partial_{i}\Pi^{i}+\mathcal{B}(\Pi^{i})^{2}+\mathcal{C}_{i}^{(0,i)}\Pi^{i}+\mathcal{D}^{(0,i)}$ and $\dot{\phi}_{1}(\vec{x}),H_{\mathrm{C}}]=\mathcal{C}_{k}^{*(0,i)}\Pi^{k}+\mathcal{D}^{*(0,i)}$, (103) implies that $\ddot{\phi}_{1}=\mathcal{C}_{k}^{*(0,i)}\Pi^{k}+\mathcal{D}^{*(0,i)}$ vanishes identically. Consequently, no additional constraints arise, and the Dirac matrix takes the form
@@
$h_{4,1}$ $=\alpha_{3},\quad f_{4,2}=\beta_{3}\quad,h_{4,3}=\gamma_{3},$ (110)
@@ VI Conclusion
We found that the complete removal of the unwanted mode requires two degeneracy conditions. The first condition is derived in Sec. IV and given in (32), which identifies the Lagrangians in (41), (42) and (46). If only the first degeneracy condition is satisfied, the theory typically propagates 2.5 DOFs which is allowed in Lorentz-violating field theories. To fully eliminate the residual half DOF, the second degeneracy condition is necessary. This is given by (50), where the Dirac matrix $B_{ab}$ is defined in (49). There are two branches of solutions to (50), denoted as the first and second branches, expressed respectively in (51) and (52).
@@
Our results show the possibility of constructing spatially covariant vector field theories that propagate only two degrees of freedom, providing a foundation for future developments in Lorentz-violating field theories and their potential cosmological applications. In this work, we considered only first-order derivatives of the vector field and focused on a flat background. It will be interesting to extend the analysis by including higher-order derivatives of the vector field and incorporating the effects of gravity. We will explore these issues in future work.
@@ Appendix A Hamiltonian formalism for Maxwell theory
The momenta conjugate to $A^{i}$ are $\Pi^{i}=\dot{A}^{i}-\partial^{i}A_{0}$. The absence of a kinetic term for the temporal component $\dot{A}_{0}$ leads to the primary constraint
$C_{1}\equiv\Pi^{0}\approx 0.$
@@
$H_{\mathrm{T}}=\int\mathrm{d}^{3}x(\mathcal{H}_{\mathrm{C}}+\lambda C_{1})$
@@
$\mathcal{H}_{\mathrm{C}}=\frac{1}{2}\Pi^{i}\Pi_{i}+\frac{1}{2}F_{ij}F^{ij}-A_{0}\partial_{i}\Pi^{i},$
@@
$\dot{C}_{1}(\vec{x})=[\Pi^{0}(\vec{x}),H_{C}]=\partial_{i}\Pi^{i}\equiv C_{2}(\vec{x})\approx 0,$
@@
$\dot{C}_{2}(\vec{x})=\left[\frac{\partial\Pi^{i}(\vec{x})}{\partial x^{i}},H_{\mathrm{C}}\right]=0.$

### v2

@@ V.2.1 Case 1
Here $\nu(\vec{y})$ is the Lagrange multiplier associated with the primary constraint $\phi_{1}$ in the total Hamiltonian, which reads $H_{\mathrm{T}}=H_{\mathrm{C}}+\int\mathrm{d}^{3}y\,\nu(\vec{y})\phi_{1}(\vec{y})$. We need to calculate the Poisson bracket $\ddot{\phi}_{1}(\vec{x}),H_{\mathrm{C}}]$ with $\ddot{\phi}_{1}=\mathcal{C}_{k}^{*(0,i)}\varPi^{k}+\mathcal{D}^{*(0,i)}$. We find
$[\ddot{\phi}_{1}(\vec{x}),H_{\mathrm{C}}]$ $=\varTheta_{m}^{(0,i)}\varPi^{m}+\varDelta^{(0,i)},$ (88)
@@
$\varTheta^{m(0,i)}=$ $\mathcal{A}\mathcal{K}\Bigg[\frac{1}{\mathcal{A}}\frac{\partial\mathcal{D}^{*(0,i)}}{\partial A_{m}}-\partial_{k}\left(\frac{\partial^{2}\mathcal{V}^{(0,i)}}{\partial A_{m}\partial A_{k}}\right)+\partial^{m}\left(\frac{\partial^{2}f_{3,2}}{\partial A_{l}\partial A_{k}}\right)\partial_{k}A_{l}$
@@
which indicates the emergence of a first-class constraint and thus the enhancement of the gauge symmetry. The resulting theory possesses one first-class constraint and two second-class constraints, and thus belongs to “type-I” theories. Comparing with the case in Sec. V.1.2, if the conditions discussed in V.1.2 are satisfied, the Poisson brackets $\phi_{1},\ddot{\phi}_{1}]\approx 0$ and $\ddot{\phi}_{1},H_{\mathrm{C}}]\approx 0$ will automatically be satisfied, but not vice versa. In other words, the case 2 of branch 1 discussed in V.1.2 is not independent in the sense that it can be viewed as the sub-case discussed here, i.e., as the special solution in case 1 of branch 2.
We then turn to the general case, where $\ddot{\phi}_{1}(\vec{x}),H_{\mathrm{C}}]$ does not vanish weakly. This leads to a new quaternary constraint, yielding the Dirac matrix
@@
$\tilde{\mathcal{F}}_{k}^{*}=-f_{4,2}\varTheta_{k}^{(0,i)}$ (94)
@@
We refer to theories in this case as “type-II” theories. It is clear that this is a new situation different from the previous ones.
@@
where $f_{2}$ is a nonlinear polynomial of $\bar{X}$. For $\mathcal{L}_{3}$, we find that $f_{3,1}$ cannot depend on $\bar{X}$ but can be a general polynomial of $A_{0}$, whereas $f_{3,2}$ must be nonlinear in $\bar{X}$. Thus we can write
@@
$h_{4,1}$ $=\alpha_{2},\quad f_{4,2}=\beta_{2},\quad h_{4,3}=\gamma_{2},$ (99)
@@ V.2.2 Case 2
Recalling that $\dot{\phi}_{1}=\mathcal{A}\partial_{i}\varPi^{i}+\mathcal{B}(\varPi^{i})^{2}+\mathcal{C}_{i}^{(0,i)}\varPi^{i}+\mathcal{D}^{(0,i)}$ and $\dot{\phi}_{1}(\vec{x}),H_{\mathrm{C}}]=\mathcal{C}_{k}^{*(0,i)}\varPi^{k}+\mathcal{D}^{*(0,i)}$, (103) implies that $\ddot{\phi}_{1}=\mathcal{C}_{k}^{*(0,i)}\varPi^{k}+\mathcal{D}^{*(0,i)}$ vanishes identically. Consequently, no additional constraints arise, and the Dirac matrix takes the form
@@
At this point, it is interesting to check the gauge transformation generated by the two first-class constraints $\phi_{1}$ and $\dot{\phi}_{1}$. Since all coefficients are constants in this case, $\dot{\phi}_{1}$ reduces to a Gauss-type constraint $\dot{\phi}_{1}=\mathcal{A}\partial_{i}\varPi^{i}\approx$ with constant $\mathcal{A}$. A convenient choice of the generator of gauge transformation is
$G[\epsilon]=\int\mathrm{d}^{3}x\left(-\dot{\epsilon}(\vec{x})\phi_{1}(\vec{x})+\frac{1}{\mathcal{A}}\epsilon(\vec{x})\dot{\phi}_{1}(\vec{x})\right),$
where $\epsilon(\vec{x})$ is the infinitesimal parameter. The transformations are thus given by
$\delta A_{0}(\vec{x})$ $=\left[A_{0}(\vec{x}),G[\epsilon]\right]=-\dot{\epsilon}(\vec{x}),$
$\delta A_{i}(\vec{x})$ $=\left[A_{i}(\vec{x}),G[\epsilon]\right]=\partial_{i}\epsilon(\vec{x}).$
This makes explicit the close analogy to electromagnetism, where the Gauss-type constraint generates $\delta A_{i}=\partial_{i}\epsilon$.
@@
$h_{4,1}$ $=\alpha_{3},\quad f_{4,2}=\beta_{3},\quad h_{4,3}=\gamma_{3},$ (110)
@@ VI Conclusion
We found that the complete removal of the unwanted mode requires two degeneracy conditions. The first condition is derived in Sec. IV and given in (32), which identifies the Lagrangians in (41), (42) and (46). If only the first degeneracy condition is satisfied, the kernel $B_{ab}(\vec{x},\vec{y})$ defined in (49) is invertible, which means all three constraints $\{\phi_{1},\dot{\phi}_{1},\ddot{\phi}_{1}\}$ are second-class. As a result, the theory typically propagates 2.5 DOFs which is allowed in Lorentz-violating field theories. To fully eliminate the residual half DOF, the second degeneracy condition is necessary. This is given by (50), where the Dirac matrix $B_{ab}$ is defined in (49). There are two branches of solutions to (50), denoted as the first and second branches, expressed respectively in (51) and (52).
@@
Our results show the possibility of constructing spatially covariant vector field theories that propagate only two degrees of freedom, providing a foundation for future developments in Lorentz-violating field theories and their potential cosmological applications. In this work, we considered only first-order derivatives of the vector field and focused on a flat background. It will be interesting to extend the analysis by including higher-order derivatives of the vector field and incorporating the effects of gravity. Moreover, a systematic extension to curved backgrounds can be carried out by promoting the building blocks to a general foliated spacetime and replacing partial derivatives by spatially covariant derivatives. We will explore these issues in future work.
@@ Appendix A Hamiltonian formalism for Maxwell theory
The momenta conjugate to $A^{i}$ are $\varPi^{i}=\dot{A}^{i}-\partial^{i}A_{0}$. The absence of a kinetic term for the temporal component $\dot{A}_{0}$ leads to the primary constraint
$C_{1}\equiv\varPi^{0}\approx 0.$
@@
$H_{\mathrm{T}}=\int\mathrm{d}^{3}x(\mathcal{H}_{\mathrm{C}}+\lambda C_{1}),$
@@
$\mathcal{H}_{\mathrm{C}}=\frac{1}{2}\varPi^{i}\varPi_{i}+\frac{1}{2}F_{ij}F^{ij}-A_{0}\partial_{i}\varPi^{i},$
@@
$\dot{C}_{1}(\vec{x})=[\varPi^{0}(\vec{x}),H_{C}]=\partial_{i}\varPi^{i}\equiv C_{2}(\vec{x})\approx 0,$
@@
$\dot{C}_{2}(\vec{x})=\left[\frac{\partial\varPi^{i}(\vec{x})}{\partial x^{i}},H_{\mathrm{C}}\right]=0.$